A wheel turns with angular speed of 30 rev s^-1 is brought to rest with constant acceleration. It turns 60 rev before it stops. Calculate its angular acceleration in rad s^-2


1
Expert's answer
2021-09-15T11:13:33-0400

By definition, the angular acceleration is given as follows:


"\varepsilon = \dfrac{\omega_f - \omega_i}{t} = \dfrac{ - \omega_i}{t}"

where "\omega_f = 0rad/s" is the final angular speed (the wheel is brought to rest), "\omega_i = 2\pi \cdot 30rev/s = 60\pi\space rad/s" is the inital angular speed, "t" is time in which it was brought to rest.

The angular distance is given as follows:


"\varphi = \omega _it + \dfrac{\varepsilon t^2}{2}"

Expressing time from the first formula and substituting in into the second one, obtain:


"t = \dfrac{ -\omega_i}{\varepsilon}\\""\varphi = \omega _i\cdot \dfrac{ -\omega_i}{\varepsilon} + \dfrac{\varepsilon\omega_i^2}{2\varepsilon^2} = -\dfrac{\omega_i^2}{2\varepsilon}"

Expressing acceleration and substituting "\varphi = 2\pi\cdot 60rev = 120\pi\space rad", obtain:


"\varepsilon = -\dfrac{\omega^2}{2\varphi} = -\dfrac{(60\pi\space rad/s)^2}{2\cdot 120\pi \space rad} = 15\pi \space rad/s^2"

Answer. "15\pi \space rad/s^2".